OpenStudy (anonymous):

can someone help me expand ln(x^2 y^4)/z^4

OpenStudy (ash2326):

Is this your question? \[\ln (\frac{x^2y^4}{z^4})\]

OpenStudy (anonymous):

Yes it it

OpenStudy (anonymous):

*is

OpenStudy (ash2326):

Do you know the addition and subtraction property of logarithms?

OpenStudy (anonymous):

no I'm not sure

OpenStudy (ash2326):

log A + log B=log AB Log A- log B= log A/B so can you use these here? don't change the power yet, just split the multiplication and division

OpenStudy (anonymous):

so i can use either the addition equation or subtraction equation ? or both?

OpenStudy (ash2326):

we can use both, but it would be good if we use step by step first split the division by using the subtraction formula

OpenStudy (anonymous):

so \[\ln x^2y^4-lnz^4\]

OpenStudy (ash2326):

good, now split the multiplication in first term

OpenStudy (ash2326):

brenda after that use this property \[\ln A^b= b\times \ln A\] Use this for each term and your question will be solved I'll be back in some time

OpenStudy (anonymous):

will it be \[2\ln x+4\ln y-3\ln z\]

OpenStudy (ash2326):

yes, good work :)

OpenStudy (ash2326):

- 4 ln z

OpenStudy (anonymous):

ohh yeah oops. anyway thank you so much for walking me through the process :)

OpenStudy (ash2326):

Welcome :)